OpenStudy (anonymous):

Algebraic structure problem:

OpenStudy (anonymous):

LeT it be G=(-1,1) with x,y from G; Show that: \[\frac{ x+y }{ 1+xy } \epsilon G\]

OpenStudy (anonymous):

@ganeshie8 @iPwnBunnies @phi @mathslover @nincompoop

OpenStudy (anonymous):

For x= y = -1 \[\frac{ x+y }{ 1+xy} = -1 \in G\] For x= y = 1 \[\frac{ x+y }{ 1+xy} = 1 \in G\] For x= 1, y = -1 \[\frac{ x+y }{ 1+xy} = 0 \in G\] Since it is true for extremes, so \[\frac{ x+y }{ 1+xy}\in G\] for x,y in (-1,1)